a small wind mill is capable of producing electrical power of 100 watt bulb to glow for four hours find the average kinetic energy of wind if the efficiency of wind mill is 14.4%

Grade:8

1 Answers

arun
8 years ago
power generated was 100 watt
now bulb can glow for 4 hours = 4\times 3600 sec
this means total energy produced by wind mill is 100\times 4\times 3600 = 0.144\times 10^7 J
now efficiency of wind mill is 14.4% = 0.144
and efficiency is ratio of energy utilised and total energy
this means total  is \frac{0.144\times 10^7}{0.144}= 10^7 J
so kinetic energy of wind is 10^7 J